A 10.0 gram sample of a compound consisting solely of sulfur and oxygen contains 4.00 grams of sulfur. What is the empirical formula of the compound?


1
Expert's answer
2021-09-21T02:00:17-0400

Molar mass of sulfur = 32.065

4/32.065 = 0.124

Molar mass of Oxygen = 15.999

6/15.999 = 0.375

0.375/0.124 = 3.02

= SO3